OpenStudy (anonymous):

Can someone please help me with this quiestion? ;I Write a paragraph proof that justifies the conditional statement “If 15 – 2x = 6 + x, then x = 3.” Use the two-column proof below to guide you.

OpenStudy (anonymous):

Statements Reasons 15 – 2x = 6 + x Given 15 = 6 + 3x Addition property of equality 9 = 3x Subtraction property of equality 3 = x Division property of equality x = 3 Symmetric property

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

But how would i put that in a paragraph?? @tgawade